These are all mixed fractions I tried to write it to the best of my ability.
The first three term and the last term is given I need to know how many terms.
3 1/8, 4 1/4, 5 3/8,..., 14 3/8

Answer by Edwin McCravy(20065)

Instead of doing your problem for you, I'll do one just like
it with mixed fractions:

2 5/16, 3 11/16, 5 1/16, ... 18 13/16

Change them to decimals with a calculator:

2.3125, 3.6875, 5.0625,.., 18.8125

common difference = d = 2nd term-1st term = 3.6875-2.3125 = 1.375 
common difference = d = 3rd term-2nd term = 5.0625-3.6875 = 1.375

      an = a1 + (n-1)d
 18.8125 = 2.3125+(n-1)1.375
 
Subtract 3.125 from both sides

    16.5 = (n-1)1.375
    16.5 = 1.375(n-1)
    16.5 = 1.375n - 1.375

Add 1.375 to both sides:
    
   17.875 = 1.1375n

Divide both sides by 1.1375:

       13 = n

Answer: 13 terms

Now do yours the exact same way.

Edwin